40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 293 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  2354   Sat Dec 5 01:40:11 2009 KojiUpdateoplevsOplevs centered, IP_POS and IP_ANG centered

We restarted daqd and it did restored the problem
http://lhocds.ligo-wa.caltech.edu:8000/40m/Computer_Restart_Procedures#fb40m
Then restart the 'daqd' process:'telnet fb40m 8087', type "shutdown" at the prompt. The framebuilder will restart itself in ~20s.

 

It did not related to the problem, but we also cleaned the processes related to dtt, dataviewer by pkill

After that the alignment scripts started to work again. As a result, we got some misalignment of the oplevs.
I am going to come on Sunday
- Align the optics
- Align the oplevs again
- Take snapshots for the suspensions
- Align the IP_POS, IP_ANG
- Align the aux laser for the absolute length
- Align PSL table QPDs, and MCT QPD

  2353   Fri Dec 4 23:17:55 2009 robUpdateoplevsOplevs centered, IP_POS and IP_ANG centered

Quote:

[Jenne Koji]

 We aligned the full IFO, and centered all of the oplevs and the IP_POS and IP_ANG QPDs.  During alignment of the oplevs, the oplev servos were disabled.

Koji updated all of the screenshots of 10 suspension screens.  I took a screenshot (attached) of the oplev screen and the QPD screen, since they don't have snapshot buttons.

We ran into some trouble while aligning the IFO.  We tried running the regular alignment scripts from the IFO_CONFIGURE screen, but the scripts kept failing, and reporting "Data Receiving Error".  We ended up aligning everything by hand, and then did some investigating of the c1lsc problem.  With our hand alignment we got TRX to a little above 1, and TRY to almost .9 . SPOB got to ~1200 in PRM mode, and REFL166Q got high while in DRM (I don't remember the number). We also saw a momentary lock of the full initerferometer:   On the camera view we saw that Yarm locked by itself momentarily, and at that same time TRX was above 0.5 - so both arms were locked simultaneously.   We accepted this alignment as "good", and aligned all of the oplevs and  QPDs.

It seems that C1LSC's front end code runs fine, and that it sees the RFM network, and the RFM sees it, but when we start running the front end code, the ethernet connection goes away.  That is, we can ping or ssh c1lsc, but once the front end code starts, those functions no longer work.  During these investigations, We once pushed the physical reset button on c1lsc, and once keyed the whole crate.  We also did a couple rounds of hitting the reset button on the DAQ_RFMnetwork screen.

 A "Data Receiving Error" usually indicates a problem with the framebuilder/testpoint manager, rather than the front-end in question.  I'd bet there's a DTT somewhere that's gone rogue.

  2352   Fri Dec 4 21:48:01 2009 JenneUpdateoplevsOplevs centered, IP_POS and IP_ANG centered

[Jenne Koji]

 We aligned the full IFO, and centered all of the oplevs and the IP_POS and IP_ANG QPDs.  During alignment of the oplevs, the oplev servos were disabled.

Koji updated all of the screenshots of 10 suspension screens.  I took a screenshot (attached) of the oplev screen and the QPD screen, since they don't have snapshot buttons.

We ran into some trouble while aligning the IFO.  We tried running the regular alignment scripts from the IFO_CONFIGURE screen, but the scripts kept failing, and reporting "Data Receiving Error".  We ended up aligning everything by hand, and then did some investigating of the c1lsc problem.  With our hand alignment we got TRX to a little above 1, and TRY to almost .9 . SPOB got to ~1200 in PRM mode, and REFL166Q got high while in DRM (I don't remember the number). We also saw a momentary lock of the full initerferometer:   On the camera view we saw that Yarm locked by itself momentarily, and at that same time TRX was above 0.5 - so both arms were locked simultaneously.   We accepted this alignment as "good", and aligned all of the oplevs and  QPDs.

It seems that C1LSC's front end code runs fine, and that it sees the RFM network, and the RFM sees it, but when we start running the front end code, the ethernet connection goes away.  That is, we can ping or ssh c1lsc, but once the front end code starts, those functions no longer work.  During these investigations, We once pushed the physical reset button on c1lsc, and once keyed the whole crate.  We also did a couple rounds of hitting the reset button on the DAQ_RFMnetwork screen.

Attachment 1: Oplev_IPang_screenshot_4Dec2009.png
Oplev_IPang_screenshot_4Dec2009.png
  2351   Fri Dec 4 18:54:03 2009 JenneUpdatePEMRanger moved

The Ranger was left in a place where it could be bumped during next week's activities (near the crawl-space to access the inside of the "L" of the IFO on the Yarm).  It has been moved a meter or so to a safer place.

Also, so that Steve can replace the battery in the SR560 that is used for the Ranger, I swapped it out with one of the ones which already has a new, charged battery.  All of the settings are identical.  For posterity, I took a pic of the front panel before unplugging the old SR560.

Attachment 1: RangerSeismometer_SR560settings_4Dec2009.JPG
RangerSeismometer_SR560settings_4Dec2009.JPG
  2350   Thu Dec 3 15:55:24 2009 AlbertoAoGLSCRF AM Stabilizer Output Power

Today I measured the max output power at the EOM output of one of the RF AM Stabilizers that we use to control the modulation depth. I needed to know that number for the designing of the new RF system.

When the EPICS slider of the 166 MHz modulation depth is at 0 the modulation depth is max (the slider's values are reversed : 0 is max, 5 is min; it is also 0 for any value above 5, sepite it range from 0 to 10).

I measured 9.5V from the EOM output, that is 32 dBm on a 50 Ohm impedance.

  2349   Mon Nov 30 19:23:50 2009 JenneUpdateMZMZ down

Came back from dinner to find the Mach Zehnder unlocked.  The poor IFO is kind of having a crappy day (computers, MZ, and I think the Mode Cleaner alignment might be bad too).

  2348   Mon Nov 30 16:23:51 2009 JenneUpdateComputersc1omc restarted

I found the FEsync light on the OMC GDS screen red.  I power cycled C1OMC, and restarted the front end code and the tpman.  I assume this is a remnant of the bootfest of the morning/weekend, and the omc just got forgotten earlier today.

  2347   Mon Nov 30 11:45:54 2009 JenneUpdateComputersWireless is back

When Alberto was parting the Red Sea this morning, and turning it green, he noticed that the wireless had gone sketchy.

When I checked it out, the ethernet light was definitely blinking, indicating that it was getting signal.  So this was not the usual case of bad cable/connector which is a known problem for our wireless (one of these days we should probably relay that ethernet cable....but not today).  After power cycling and replugging the ethernet cable, the light for the 2.4GHz wireless was blinking, but the 5GHz wasn't.  Since the wireless still wasn't working, I checked the advanced configuration settings, as described by Yoichi's wiki page:  40m Network Page

The settings had the 5GHz disabled, while Yoichi's screenshots of his settings showed it enabled.  Immediately after enabling the 5GHz, I was able to use the laptop at Alberto's length measurement setup to get online.  I don't know how the 5GHz got disabled, unless that happened during the power cycle (which I doubt, since no other settings were lost), but it's all better now.

 

  2346   Mon Nov 30 11:29:40 2009 AlbertoAoGall down cond.sea of red

Quote:

Quote:

Came in, found all front-ends down.

 

Keyed a bunch of crates, no luck:

Requesting coeff update at 0x40f220 w/size of 0x1e44
No response from EPICS 

Powered off/restarted c1dcuepics.  Still no luck.

Powered off megatron.  Success!  Ok, maybe it wasn't megatron.  I also did c1susvme1 and c1susvme2 at this time.

 

BURT restored to Nov 26, 8:00am

 

But everything is still red on the C0_DAQ_RFMNETWORK.adl screen, even though the front-ends are running and synced with the LSC.  I think this means the framebuilder or the DAQ controller is the one in trouble--I keyed the crates with DAQCTRL and DAQAWG a couple of times, with no luck, so it's probably fb40m.    I'm leaving it this way--we can deal with it tomorrow.

 I found the red sea when I came in this morning.

I tried several things.

  1. ssh into fb40m: connection refused
  2. telnet fb40m 8087: didn't respond
  3. shutdown fb40m by physically pushing the power button: it worked and the FB came back to life but still with a red light on the MEDM DAQ_DETAIL screen;
  4. powercycled fb40m AND C0DAQCTRL: no improvement
  5. shutdown fb40m, C0DAQCTRL, C1DCUEPICS and pushed the reset button on the RF network crate; then I restarted the computers in this order: fb40m, C1DCUEPICS, C0DAQCTRL: it worked: they came back to life and the lights eventually turned green on the MEDM montior screen

I'm now going to restart the single front -ends and burtgooey them if necessary.

Everything is back on.

Restarted all the front ends. As usual c1susvme2 was stubborn  but eventually it came up.

I burt-restored all the front-ends to Nov 26 at 8am.

The mode cleaner is locked.

  2345   Mon Nov 30 10:28:47 2009 AlbertoAoGall down cond.sea of red

Quote:

Came in, found all front-ends down.

 

Keyed a bunch of crates, no luck:

Requesting coeff update at 0x40f220 w/size of 0x1e44
No response from EPICS 

Powered off/restarted c1dcuepics.  Still no luck.

Powered off megatron.  Success!  Ok, maybe it wasn't megatron.  I also did c1susvme1 and c1susvme2 at this time.

 

BURT restored to Nov 26, 8:00am

 

But everything is still red on the C0_DAQ_RFMNETWORK.adl screen, even though the front-ends are running and synced with the LSC.  I think this means the framebuilder or the DAQ controller is the one in trouble--I keyed the crates with DAQCTRL and DAQAWG a couple of times, with no luck, so it's probably fb40m.    I'm leaving it this way--we can deal with it tomorrow.

 I found the red sea when I came in this morning.

I tried several things.

  1. ssh into fb40m: connection refused
  2. telnet fb40m 8087: didn't respond
  3. shutdown fb40m by physically pushing the power button: it worked and the FB came back to life but still with a red light on the MEDM DAQ_DETAIL screen;
  4. powercycled fb40m AND C0DAQCTRL: no improvement
  5. shutdown fb40m, C0DAQCTRL, C1DCUEPICS and pushed the reset button on the RF network crate; then I restarted the computers in this order: fb40m, C1DCUEPICS, C0DAQCTRL: it worked: they came back to life and the lights eventually turned green on the MEDM montior screen

I'm now going to restart the single front -ends and burtgooey them if necessary.

  2344   Sun Nov 29 16:56:56 2009 robAoGall down cond.sea of red

Came in, found all front-ends down.

 

Keyed a bunch of crates, no luck:

Requesting coeff update at 0x40f220 w/size of 0x1e44
No response from EPICS 

Powered off/restarted c1dcuepics.  Still no luck.

Powered off megatron.  Success!  Ok, maybe it wasn't megatron.  I also did c1susvme1 and c1susvme2 at this time.

 

BURT restored to Nov 26, 8:00am

 

But everything is still red on the C0_DAQ_RFMNETWORK.adl screen, even though the front-ends are running and synced with the LSC.  I think this means the framebuilder or the DAQ controller is the one in trouble--I keyed the crates with DAQCTRL and DAQAWG a couple of times, with no luck, so it's probably fb40m.    I'm leaving it this way--we can deal with it tomorrow.

  2343   Sat Nov 28 20:27:12 2009 KojiUpdatePSLFSS oscillation: Total gain reduced

I stopped by the 40m for some reason and found that the MC trans was 7.5.
This was caused by an oscillation of FSS, which seemed to be started by itself.

The oscillation stopped by reducing the FSS total gain to +9dB (from +11dB).
This is not a permanent fix (i.e. autolocker will restore the gain).
If it seems necessary to reduce the FSS gain always, we change the MC autolocker script.

Attachment 1: 091128_PSL.png
091128_PSL.png
  2342   Fri Nov 27 02:25:26 2009 ranaUpdateABSLPLL Open Loop Gain Measured

Quote:

I measured the open loop gain of the PLL in the AbsL experiment.

 Plots don't really make sense. The second one is inherently unstable - and what's g?

  2341   Thu Nov 26 02:08:34 2009 KojiUpdateElectronicsMulti-resonant EOM --- Q-factor ----

The key point of the story is:
"The recipe to exploit maximum benefit from a resonant EOM"
- Make a resonant EOM circuit. Measure the impedance Z at the resonance.
- This Z determines the optimum turn ratio n of the step-up transformer.
 
(n2 = Z/Rin where Rin is 50Ohm in our case.)
- This n gives the maximum gain Gmax (= n/2) that can be obtained with the step up transformer.
  And, the impedance matching is also satisfied in this condition.

OK: The larger Z, the better. The higher Q, the Z larger, thus the better.
(Although the relationship between Z and Q were not described in the original post.)

So, how can we make the Q higher? What is the recipe for the resonant circuit?
=> Choose the components with smaller loss (resistance). The details will be provided by Kiwamu soon??? 


When I was young (3 months ago), I thought...

  • Hey! Let's increase the Q of an EOM! It will increase the modulation!
  • Hey! Let's use the step-up transformer with n as high as possible! It will increase the modulation!
  • Hey! Take the impedance matching! It will increase the modulation!

I was just too thoughtless. In reality, they are closely related each other.

A high Q resonant circuit has a high residual resistance at the resonant frequency. As far as the impedance is higher than the equivalent output impedance of the driving circuit (i.e. Z>Rin n2), we get the benefit of increasing the turn ratio of the transformer. In other words, "the performance of the resonant EOM is limited by the turn ratio of the transformer." (give us more turns!)

OK. So can we increase the turn ratio infinitely? No. Once Rin n2 gets larger than Z, you no longer get the benefit of the impedance transforming. The output impedance of the signal source yields too much voltage drop.

There is an optimum point for n. That is the above recipe. 

So, a low Q resonant EOM has a destiny to be useless. But high Q EOM still needs to be optimized. As far as we use a transformer with a low turn ratio, it only shows ordinary performance.

 

 

  2340   Wed Nov 25 20:44:48 2009 kiwamuUpdateElectronicsMulti-resonant EOM --- Q-factor ----

Now I am studying about the behavior of the Q-factor in the resonant circuit because the Q-factor of the circuit directly determine the performance as the EOM driver.

Here I summarize the fundamental which explains why Q-factor is important.

 --------------------------------------

The EOM driver circuit can be approximately described as shown in figure below

trans.png

Z represents the impedance of a resonant circuit.

In an ideal case, the transformer just raise the voltage level n-times larger.  Rin is the output impedance of the signal source and usually has 50[Ohm].

The transformer also makes the impedance Z 1/n^2 smaller. Therefore this configuration gives a following relation between Vin and Vout.

eq1.png

 Where G is the gain for the voltage. And G goes to a maximum value when Rin=Z/n2. This relation is shown clearly in the following plot.

 

impedance.png

 Note that I put Rin=50 [Ohm] for calculating the plot.

Under the condition  Rin=Z/n2( generally referred as impedance matching ), the maximum gain can be expressed as;

eq2.png

 

It means that larger Z makes more efficient gain. In our case, interested Z is considered as the impedance at a resonance.

So what we should do is making a resonant circuit which has a higher impedance at the resonance (e.g. high Q-resonant circuit).

 

 

  2339   Wed Nov 25 20:28:17 2009 AlbertoUpdateABSLStopped working on the AbsL

I closed the shutter of the NPRO for the night.

  2338   Wed Nov 25 20:24:49 2009 AlbertoUpdateABSLPLL Open Loop Gain Measured

I measured the open loop gain of the PLL in the AbsL experiment.

I repeated the measurement twice: one with gain knob on the universal PDH box g=3.0; the second measurement with g=6.0

The UGF were 60 KHz and 100 KHz, respectively.

That means that one turn of the knob equals to about +10 dB.

Attachment 1: 2009-09-25_OLgain_g3png.png
2009-09-25_OLgain_g3png.png
Attachment 2: 2009-09-25_OLgain_g6png.png
2009-09-25_OLgain_g6png.png
  2337   Wed Nov 25 20:14:58 2009 AlbertoUpdateABSLAbsL PLL not able to lock: problem fixed

Quote:

Last night something happened on the beat between the PSL beam and the auxiliary NPRO beam, that spoiled the quality of the beating I had before. As a result the PLL has become unable to lock the two lasers.

The amplitude of the beat at the spectrum analyzer has gone down to -40 dBm from -10 that it was earlier. The frequency has also become more unstable so that now it can be seen writhing within tens of KHz.

Meanwhile the power of the single beams at the PLL photodiode hasn't changed, suggesting that the alignment of the two beam didn't change much.

Changes in the efficiency of the beating between the two beams are not unusual. Although that typically affects only the amplitude of the beat and wouldn't explain why also its frequency has become unstable. Tuning the alignment of the PLL optics usually brings the amplitude back, but it was uneffective today.

It looks like something changed in either one of the two beams. In particular the frequency of one of the two lasers has become less stable.

Another strange thing that I've been observing is that the amplitude of the beat goes down (several dBm) as the beat frequency is pushed below 50 MHz. Under 10 MHz it even gets to about -60 dBm.

I noticed the change yesterday evening at about 6pm, while I was taking measurements of the PLL open loop tranfer function and everything was fine. I don't know whether it is just a coincidence or it is somehow related to this, but Jenne and Sanjit had then just rebooted the frame builder.

 

Problem found. Inspecting with Koji we found that there was a broken SMA-to-BNC connector in the BNC cable from the photodiode.

  2336   Wed Nov 25 16:44:52 2009 KojiUpdatePSLMeasured MC length--FSS trend

I checked C1:PSL-FSS_VCODETPWR. The attached is the 4 months trend of the FSS RCTRANS / RFPDDC(=FSS REFL) / VCODETPWR / VCOMODLEVEL.

Although VCO modulation level setting was mostly constnt, VCODETPWR, which presumably represents the RF level, changes time by time.
It coincides with the recent reduction of the RCTRANS/RFPDDC. Actually, my touch restored the VCO to the previous more stable state.
One can see that this is not only a single occation, but it happened before too. (In the middle of Aug.)

This could be explained by the bad contact of some cable or connector.

Nevertheless we need more careful investigation:

1. Understand what VCODETPWR is exactly.
2. Investigate relationship between VCOMODLEVEL / VCODETPWR / AOM deflection efficiency / RCTRANSPD
3. Confirm the frequency matching between the VCO and AOM.

Quote:

but the increase in both the RCtrans and the RCrefl is consistent with my theory that the power going to the RC has increased ; its not just an increase in the visibility.

We should scan the AOM/VCO to make sure the frequency is matched to the resonance to within 0.5 dB.

 

Attachment 1: 091125_FSS.png
091125_FSS.png
  2335   Wed Nov 25 16:13:27 2009 ranaUpdatePSLMeasured MC length--FSS trend

but the increase in both the RCtrans and the RCrefl is consistent with my theory that the power going to the RC has increased ; its not just an increase in the visibility.

We should scan the AOM/VCO to make sure the frequency is matched to the resonance to within 0.5 dB.

  2334   Wed Nov 25 15:42:27 2009 AlbertoUpdateABSLWorking on the AP table

Quote:

I'm working on the AP table. I also opened the auxiliary NPRO shutter. The auxiliary beam is on its path on the AP table and PSL table.

 NPRO shutter closed

  2333   Wed Nov 25 15:38:08 2009 robUpdateLockingMeasured MC length

Quote:

Quote:

What I meant was the VCO driver, not the FSS box.

As for the frequency, all written numbers were the Marconi displays.
The number on the frequency counter was also recorded, and so will be added to the previous entry shortly... 

Quote:

I propose that from now on, we indicate in the elog what frequencies we're referring to. In this case, I guess its the front panel readback and not the frequency counter -- what is the frequency counter readback? And is everything still locked to the 10 MHz from the GPS locked Rubidium clock?

Plus, what FSS Box? The TTFSS servo box? Or the VCO driver? As far as I know, the RC trans PD doesn't go through the FSS boxes, and so its a real change. I guess that a bad contact in the FSS could have made a huge locking offset.

 

 

Locking has gone sour.  The CARM to MCL handoff, which is fairly early in the full procedure and usally robust, is failing reliably. 

As soon as the SUS-MC2_MCL gain is reduced, lock is broken.  There appears to be an instability around 10Hz.  Not sure if it's related.

 Whatever the locking problem was, the power of magical thinking has forced it to retreat for now.  The IFO is currently locked, having completed the full up script.  One more thing for which to be thankful.

  2332   Wed Nov 25 14:29:08 2009 robUpdateLockingMeasured MC length--FSS trend

Quote:

Quote:

What I meant was the VCO driver, not the FSS box.

As for the frequency, all written numbers were the Marconi displays.
The number on the frequency counter was also recorded, and so will be added to the previous entry shortly... 

Quote:

I propose that from now on, we indicate in the elog what frequencies we're referring to. In this case, I guess its the front panel readback and not the frequency counter -- what is the frequency counter readback? And is everything still locked to the 10 MHz from the GPS locked Rubidium clock?

Plus, what FSS Box? The TTFSS servo box? Or the VCO driver? As far as I know, the RC trans PD doesn't go through the FSS boxes, and so its a real change. I guess that a bad contact in the FSS could have made a huge locking offset.

 

 

Locking has gone sour.  The CARM to MCL handoff, which is fairly early in the full procedure and usally robust, is failing reliably. 

As soon as the SUS-MC2_MCL gain is reduced, lock is broken.  There appears to be an instability around 10Hz.  Not sure if it's related.

 Five day minute trend.  FAST_F doesn't appear to have gone crazy.

Attachment 1: FSStrendpowerjump.png
FSStrendpowerjump.png
  2331   Wed Nov 25 12:28:22 2009 JenneUpdateSUSMC2 tripped

Just felt a big "kerplunk" type ground-shaking, presumably from all the antics next door.  MC2's watchdog tripped as a result.  The watchdog has been reenabled.

  2330   Wed Nov 25 11:10:05 2009 JenneUpdateComputers40m frame builder backup acting funny

Quote:

Quote:

As part of the fb40m restart procedure (Sanjit and I were restarting it to add some new channels so they can be read by the OAF model), I checked up on how the backup has been going.  Unfortunately the answer is: not well.

I'll check in with the frame builder again tomorrow, to make sure that it's all good.

 All is well again in the world of backups.  We are now up to date as of ~midnight last night. 

 Backup Fail.  At least this time however, it threw the appropriate error code, and sent me an email saying that it was unhappy.  Alan said he was going to check in with Stuart regarding the confusion with the ssh-agent.  (The other day, when I did a ps -ef | grep agent, there were ~5 ssh-agents running, which could have been then cause of the unsuccessful backups without telling me that they failed.  The main symptom is that when I first restart all of the ssh-agent stuff, according to the directions in the Restart fb40m Procedures, I can do a test ssh over to ldas-cit, to see what frames are there.  If I log out of the frame builder and log back in, then I can no longer ssh to ldas-cit without a password.  This shouldn't happen....the ssh-agent is supposed to authenticate the connection so no passwords are necessary.) 

I'm going to restart the backup script again, and we'll see how it goes over the long weekend. 

  2329   Wed Nov 25 11:02:54 2009 AlbertoUpdateABSLAbsL PLL not able to lock

Quote:

Last night something happened on the beat between the PSL beam and the auxiliary NPRO beam, that spoiled the quality of the beating I had before. As a result the PLL has become unable to lock the two lasers.

The amplitude of the beat at the spectrum analyzer has gone down to -40 dBm from -10 that it was earlier. The frequency has also become more unstable so that now it can be seen writhing within tens of KHz.

Meanwhile the power of the single beams at the PLL photodiode hasn't changed, suggesting that the alignment of the two beam didn't change much.

Changes in the efficiency of the beating between the two beams are not unusual. Although that typically affects only the amplitude of the beat and wouldn't explain why also its frequency has become unstable. Tuning the alignment of the PLL optics usually brings the amplitude back, but it was uneffective today.

It looks like something changed in either one of the two beams. In particular the frequency of one of the two lasers has become less stable.

Another strange thing that I've been observing is that the amplitude of the beat goes down (several dBm) as the beat frequency is pushed below 50 MHz. Under 10 MHz it even gets to about -60 dBm.

I noticed the change yesterday evening at about 6pm, while I was taking measurements of the PLL open loop tranfer function and everything was fine. I don't know whether it is just a coincidence or it is somehow related to this, but Jenne and Sanjit had then just rebooted the frame builder.

 I confirm what I said earlier. The amplitude of the beat is -10 dBm at 300MHz. It goes down at lower frequencies. In particular it gets to-60 dBm below 20 MHz. For some strange reason that I couldn't explain the beating efficiency has become poorer at low frequencies.

  2328   Wed Nov 25 10:20:47 2009 AlbertoUpdateABSLAbsL PLL not able to lock

Last night something happened on the beat between the PSL beam and the auxiliary NPRO beam, that spoiled the quality of the beating I had before. As a result the PLL has become unable to lock the two lasers.

The amplitude of the beat at the spectrum analyzer has gone down to -40 dBm from -10 that it was earlier. The frequency has also become more unstable so that now it can be seen writhing within tens of KHz.

Meanwhile the power of the single beams at the PLL photodiode hasn't changed, suggesting that the alignment of the two beam didn't change much.

Changes in the efficiency of the beating between the two beams are not unusual. Although that typically affects only the amplitude of the beat and wouldn't explain why also its frequency has become unstable. Tuning the alignment of the PLL optics usually brings the amplitude back, but it was uneffective today.

It looks like something changed in either one of the two beams. In particular the frequency of one of the two lasers has become less stable.

Another strange thing that I've been observing is that the amplitude of the beat goes down (several dBm) as the beat frequency is pushed below 50 MHz. Under 10 MHz it even gets to about -60 dBm.

I noticed the change yesterday evening at about 6pm, while I was taking measurements of the PLL open loop tranfer function and everything was fine. I don't know whether it is just a coincidence or it is somehow related to this, but Jenne and Sanjit had then just rebooted the frame builder.

  2326   Wed Nov 25 08:43:08 2009 AlbertoUpdateABSLWorking on the AP table

I'm working on the AP table. I also opened the auxiliary NPRO shutter. The auxiliary beam is on its path on the AP table and PSL table.

  2325   Wed Nov 25 03:05:15 2009 robUpdateLockingMeasured MC length

Quote:

What I meant was the VCO driver, not the FSS box.

As for the frequency, all written numbers were the Marconi displays.
The number on the frequency counter was also recorded, and so will be added to the previous entry shortly... 

Quote:

I propose that from now on, we indicate in the elog what frequencies we're referring to. In this case, I guess its the front panel readback and not the frequency counter -- what is the frequency counter readback? And is everything still locked to the 10 MHz from the GPS locked Rubidium clock?

Plus, what FSS Box? The TTFSS servo box? Or the VCO driver? As far as I know, the RC trans PD doesn't go through the FSS boxes, and so its a real change. I guess that a bad contact in the FSS could have made a huge locking offset.

 

 

Locking has gone sour.  The CARM to MCL handoff, which is fairly early in the full procedure and usally robust, is failing reliably. 

As soon as the SUS-MC2_MCL gain is reduced, lock is broken.  There appears to be an instability around 10Hz.  Not sure if it's related.

  2324   Tue Nov 24 19:16:02 2009 AlbertoUpdateABSLworking on the AP table

Quote:

I'm working on the AP table. I also opened the auxiliary NPRO shutter. The auxiliary beam is on its path on the AP table and PSL table.

 Closing the AP table and the NPRO shutter now.

  2323   Tue Nov 24 18:24:54 2009 SanjitConfigurationAdaptive FilteringASS channels added to framebuilder

 

[Sanjit, Jenne, Rob, Joe]

 

We added and tested the following channels from "/cvs/cds/gds/param/tpchn_C3.par" to "/cvs/cds/caltech/chans/daq/C1ASS.ini" appending a "_2048" extension to the channel name (as the name of a channel in .ini and .par files must be different):

[C1:ASS-TOP_CORR_IN1_2048]
[C1:ASS-TOP_ERR_EMPH_IN1_2048]
[C1:ASS-TOP_PEM_10_IN1_2048]
[C1:ASS-TOP_PEM_11_IN1_2048]
[C1:ASS-TOP_PEM_12_IN1_2048]
[C1:ASS-TOP_PEM_15_IN1_2048]
[C1:ASS-TOP_PEM_16_IN1_2048]
[C1:ASS-TOP_PEM_17_IN1_2048]
[C1:ASS-TOP_PEM_18_IN1_2048]
[C1:ASS-TOP_PEM_19_IN1_2048]
[C1:ASS-TOP_PEM_20_IN1_2048]
[C1:ASS-TOP_PEM_24_IN1_2048]
[C1:ASS-TOP_PEM_2_IN1_2048]
[C1:ASS-TOP_PEM_3_IN1_2048]
[C1:ASS-TOP_PEM_4_IN1_2048]
 

These five-line entries for each channels in the .par file were manually copy pasted from the .ini file, should think about a smarter way...

The old .par file is kept as: /cvs/cds/caltech/chans/daq/C1ASS.ini.20Nov2009

The current one is also saved as: /cvs/cds/caltech/chans/daq/C1ASS.ini.24Nov2009

And, the current one is committed to the svn.

 

NOTE: In the first attempt, the channel names were mistakenly kept the same in both the .ini and .par files and this caused DAQ daemon to crash badly. It could only be recovered by hard reboot of the frame builder.  Important info here: Jenne's elog 2316

  2322   Tue Nov 24 16:06:45 2009 JenneUpdateComputers40m frame builder backup acting funny

Quote:

As part of the fb40m restart procedure (Sanjit and I were restarting it to add some new channels so they can be read by the OAF model), I checked up on how the backup has been going.  Unfortunately the answer is: not well.

I'll check in with the frame builder again tomorrow, to make sure that it's all good.

 All is well again in the world of backups.  We are now up to date as of ~midnight last night. 

  2321   Tue Nov 24 14:33:22 2009 AlbertoUpdateABSLworking on the AP table

I'm working on the AP table. I also opened the auxiliary NPRO shutter. The auxiliary beam is on its path on the AP table and PSL table.

  2320   Tue Nov 24 10:36:21 2009 KojiUpdateLSCMeasured MC length

What I meant was the VCO driver, not the FSS box.

As for the frequency, all written numbers were the Marconi displays.
The number on the frequency counter was also recorded, and so will be added to the previous entry shortly... 

Quote:

I propose that from now on, we indicate in the elog what frequencies we're referring to. In this case, I guess its the front panel readback and not the frequency counter -- what is the frequency counter readback? And is everything still locked to the 10 MHz from the GPS locked Rubidium clock?

Plus, what FSS Box? The TTFSS servo box? Or the VCO driver? As far as I know, the RC trans PD doesn't go through the FSS boxes, and so its a real change. I guess that a bad contact in the FSS could have made a huge locking offset.

 

  2319   Tue Nov 24 08:00:16 2009 ranaUpdateLSCMeasured MC length

I propose that from now on, we indicate in the elog what frequencies we're referring to. In this case, I guess its the front panel readback and not the frequency counter -- what is the frequency counter readback? And is everything still locked to the 10 MHz from the GPS locked Rubidium clock?

Plus, what FSS Box? The TTFSS servo box? Or the VCO driver? As far as I know, the RC trans PD doesn't go through the FSS boxes, and so its a real change. I guess that a bad contact in the FSS could have made a huge locking offset.

 

  2318   Mon Nov 23 21:36:38 2009 KojiUpdateIOOAligned PMC/RC

I aligned the beam goes to PMC. It increased the MC Trans from 8.25 to 8.30.

I also aligned the beam goes to RC.
When I touched the FSS box (wrong: this was the VCO driver) that was close to one of the steering mirror, suddenly the RC trans increased.
It is now 9.8. I am afraid that it gets saturated. I could not reproduce the phenomenon. This could be caused by a bad contact?
Note that I didn't see there is any loose optic.

Attachment 1: 091123_PSL.png
091123_PSL.png
  2317   Mon Nov 23 21:30:29 2009 JenneUpdateLSCMeasured MC length

With Koji's help, I measured the length of the Mode Cleaner.

The new modulation frequencies (as quoted on the Marconi front panels) are: 

165.980580 MHz

 33.196116 MHz

132.784464 MHz

199.176696 MHz

The Frequency Counter readback is 165980584.101 Hz (a 4Hz difference).  All of the Marconi's front-panel frequencies read ###.##### MHz Ext, and the Frequency standard has it's "locked" light illuminated, and the 1pps input light blinking, so I think everything is still nicely locked to the frequency standard, and the frequency standard is locked to the GPS.

While changing the marconi's, I accidentally touched the MC's 29.5 MHz marconi.  It is set back to the nominal value (according to Kiwamu's rack photos) of 29.485MHz.  But the phase might be sketchy, although hopefully this doesn't matter since we don't do a double demodulation with it.

I also ran the scripts in the wiki page: How To/Diagonalize DRMI Length Control to set the DD Phases.

 

 

  2316   Mon Nov 23 19:36:28 2009 JenneUpdateAdaptive FilteringHow to add ASS channels, so that they're saved to frames

[Jenne, Sanjit]

We would like several channels from the OAF/ASS screen to be saved to frames, so that we can use the channels for our OAF model.  In theory, this should involve uncommenting the desired channels in the .ini file (.../caltech/chans/daq/C1ASS.ini), and restart the frame builder.  Since this .ini file was generated a long time ago, and things have been changed since then, the chnnums in the .ini file and the corresponding .par file don't match up.  We need to go through the .par file (/cvs/cds/gds/param/tpchn_c3.par), and look up the chnnums for our channels, and copy those numbers into the .ini file.  Figuring out what was going on involved many fb40m restarts, but on the last one of the night, I restarted the backup script, so it should (hopefully) run tonight, and get all of the frames that we've been missing.

Notes to self: 

*  When adding channels to other front ends, the end of the process is to click the blue button on the C0DAQ_DETAIL screen next to your computer.  C1ASS isn't on that screen.  Instead, in the C1ASS_GDS screen, click DAQ Reload.

*  The channel names for the Test Points and the .ini files must be different.  That's why there's a '_2048' suffix at the end of every channel in our .ini file.

*  tpchn_C1 is all of the old-style system test points.  tpchn_C2 is the C1OMC, and tpchn_C3 is for the C1ASS testpoints.

*  When uncommenting channels in the C1ASS.ini file, make sure acquire is set to 1 for every channel we want saved.  The default in this .ini file is set to acquire = 0.

  2315   Mon Nov 23 17:53:08 2009 JenneUpdateComputers40m frame builder backup acting funny

As part of the fb40m restart procedure (Sanjit and I were restarting it to add some new channels so they can be read by the OAF model), I checked up on how the backup has been going.  Unfortunately the answer is: not well.

Alan imparted to me all the wisdom of frame builder backups on September 28th of this year.  Except for the first 2 days of something having gone wrong (which was fixed at that time), the backup script hasn't thrown any errors, and thus hasn't sent any whiny emails to me.  This is seen by opening up /caltech/scripts/backup/rsync.backup.cumlog , and noticing that  after October 1, 2009, all of the 'errorcodes' have been zero, i.e. no error (as opposed to 'errorcode 2' when the backup fails).  

However, when you ssh to the backup server to see what .gwf files exist, the last one is at gps time 941803200, which is Nov 9 2009, 11:59:45 UTC.  So, I'm not sure why no errors have been thrown, but also no backups have happened. Looking at the rsync.backup.log file, it says 'Host Key Verification Failed'.  This seems like something which isn't changing the errcode, but should be, so that it can send me an email when things aren't up to snuff.  On Nov 10th (the first day the backup didn't do any backing-up), there was a lot of Megatron action, and some adding of StochMon channels.  If the fb was restarted for either of these things, and the backup script wasn't started, then it should have had an error, and sent me an email.  Since any time the frame builder's backup script hasn't been started properly it should send an email, I'm going to go ahead and blame whoever wrote the scripts, rather than the Joe/Pete/Alberto team.

Since our new raid disk is ~28 days of local storage, we won't have lost anything on the backup server as long as the backup works tonight (or sometime in the next few days), because the backup is an rsync, so it copies anything which it hasn't already copied.  Since the fb got restarted just now, hopefully whatever funny business (maybe with the .agent files???) will be gone, and the backup will work properly. 

I'll check in with the frame builder again tomorrow, to make sure that it's all good.

  2314   Mon Nov 23 16:28:12 2009 steveSummaryCamerasVideo swicher options

Quote:

Steve is summarizing the Video Matrix choices into this Wiki page:

http://lhocds.ligo-wa.caltech.edu:8000/40m/Electronics/VideoMUX

Requirements:

Price: < 5k$

Control: RS-232 and Ethernet

Interface: BNC (Composite Video)

Please check into the page on Monday for a final list of choices and add comments to the wiki page.

 Composite video matrix switchers with 32 BNC in and 32 BNC channels out are listed.

  2313   Mon Nov 23 11:03:00 2009 steveUpdateSUSjackhammer special well under control

Quote:

I've changed the watchdog rampdown script so it brings the SUS watchdogs to 220, instead of the 150 it previously targeted.  This is to make tripping less likely with the jackhammering going on next door.  I've also turned off all the oplev damping.

 Saturday's special event of braking up the large concrete pieces in CES bay was un event full.

Attachment 1: dig10d.png
dig10d.png
Attachment 2: P1050740.JPG
P1050740.JPG
  2312   Mon Nov 23 10:11:03 2009 steveUpdatePEMlong term temp fluctuation of the 40m lab

Quote:

This first plot shows the RC temperature channels' performance from 40 days ago, before we disabled the MINCO PID controller. Although RCTEMP is supposed to be the out of loop sensor, what we really care about is the cavity length and so I've plotted the SLOW. To get the SLOW on the same scale, I've multiplied the channel by 10 and then adjusted the offset to get it on the same scale.

 The second plot shows a period after that where there is no temperature control of the can at all. Same gain scaling has been applied to SLOW as above, so that instead of the usual 1 GHz/V this plot shows it in 0.1 GHz/V.

The third plot shows it after the new PID was setup.

Summary: Even though the PID loop has more gain, the true limit to the daily fluctuations in the cavity temperature and the laser frequency are due to the in-loop sensors measuring the wrong thing. i.e. the out-of-loop temperature is too different from the in-loop sensor. This can possibly be cured with better foam and better placement of the temperature sensors. Its possible that we're now just limited by the temperature gradients on the can.

 Here is a 7 years plot of  of the 40m temperature variations.

Attachment 1: 7ytemp.jpg
7ytemp.jpg
  2311   Mon Nov 23 00:46:09 2009 rana, robUpdatePSLISS RIN: Its too high by 10x

This plot shows the RIN as measured by the ISS. Its ~2 x 10^-7, whereas its supposed to be more like 3 x 10^-8.

The ISS has DC coupled RIN channels (with a _F suffix) and AC coupled RIN channels (with a _FW suffix). By using a swept sine, Rob determined that the AC coupled channels have an AC coupling pole at ~80 Hz. The attached plot uses this and then has the overall gain adjusted to match with the _F channels below 10 Hz.

The _F channels can be converted directly into RIN by just dividing the spectra by the mean value of the time series. The dark offset of these channels is small and so this only introduces a ~5-10% calibration error.

Question #1: Why is the RIN so bad? According to the MEDM screen, the photocurrent on the MON/SENS PDs is 1.9/1.3 mA. That's sort of low, but should still allow us to get 5x10^-8 in RIN.

Question #2: Does it make an effect on the current DC Readout work? IF so, should we try to fix up the ISS in a temporary way? Since the in-loop and out-of-loop detectors are completely coherent, all of the noise is likely just unsuppressed noise from the laser. We are unable to increase the gain because of the high frequency noise from the NPRO.

 

Let's remember to replace this ISS with a new one that can drive an AOM. Need a volunteer to get us a new ISS.

 

Attachment 1: Untitled.png
Untitled.png
  2310   Fri Nov 20 17:44:38 2009 JenneUpdateAdaptive FilteringSome svn shenanigans

[Sanjit, Jenne]

Sanjit and I are trying to put names to some signals which exist in SimuLink land, but which don't (yet) exist in EPICS land.  The deelio is that for each of the chosen SEIS signals in the ASS_TOP_PEM screen, the signal is split.  One part of the signal is used to decide how the adaptive filter should look, and the other part is actually used when doing the on-line subtraction.  Previously only the part of the signal which is used to decide on the Adaptive Filter could be seen on the screens, and had names. 

Before touching anything on the Simulink ASS.mdl, I did an svn check in, which put things at revision 36639. 

To try to make the desired signals exist, I put cdsFilt boxes (to create filter modules for each of these signals), and gave each of them a name (kind of like the Neverending Story....once they have a name, they'll exist).  My new names are C1:ASS-TOP_PEM_#_APPLY, which correspond to the previously-existing C1:ASS-TOP_PEM_#_ADPT (these are the ones that are along the top of the ASS_TOP_PEM matrix screen).  This version of the simulink model was checked in, and the svn is now at revision 36640.

We then did some "make clean", "make ass" and "make install-ass" action, and burt restored c1assepics, but nothing seems to be happening.  The screen doesn't have white boxes all over the place, and we didn't get any errors when we did the makes, and I'm sure we burt restored correctly (made sure the ASS GDS screen had a 1 in the lower left box etc), but all the values on the screen are still zero.  

When we ran the ass front end in terminal on the c1ass machine, we did see an error: "Invalid chan num found 2 = 30624" "DAQ init failed -- exiting".  I think this means that we need to have told some file somewhere that I was going to be adding 8 new channels. (maybe an .ini file?) Hopefully the Joe & Peter team can help us out with this, since they've been doing this kind of thing for the new system.

Moral of the story is, the new (non-working) simulink file has been svn checked in as revision 36640, and we're reverting to revision 36639, which was before I touched anything today.

  2309   Fri Nov 20 16:18:56 2009 robConfigurationSUSwatchdog rampdown

I've changed the watchdog rampdown script so it brings the SUS watchdogs to 220, instead of the 150 it previously targeted.  This is to make tripping less likely with the jackhammering going on next door.  I've also turned off all the oplev damping.

  2308   Fri Nov 20 13:54:45 2009 AlbertoOmnistructureEnvironmentNew Paper and Cardboard Recycling Bins Introduced to the 40m

The 40m produces a large amount of paper and cardboard waste. It would be worth it if we disposed those kind of garbages for recycling.

I set up two new garbage bins in the lab: one in the control rooms that adds up to the can/bottle recycling bin, and one other one in the office area, next to the printer for general paper disposal.

People in the lab are strongly invited to make use of the two new garbage bins and recycle their paper and cardboard waste.

Boxes can be larger than the garbage bin, so I'd recommend people to crash them before throwing them into the bin.

DSC_1030.JPG

DSC_1031.JPG

  2307   Fri Nov 20 11:32:48 2009 HaixingUpdateSUSMagnetic levitation

I added an integrator to increase the gain at low frequencies (below 5 Hz). In addition, I increased

the band of the differentiator. The schematics for both integrator and differentiator are the following:

IntDiff.PNG

The magnetic is stably levitated.

S8007340.jpg

I turned off the light to get rid of 60Hz noise on the photodiode. I tried to measured the

open-loop transfer function of this setup, but somehow the SR560 is always saturate

when I injected the signal from SR785, which produces some weird results at

low-frequencies.

 

In addition, I found out that when the light is turned on, the levitation

can be stable even when I inverted the sign of the control loop. The control signal

on the osciloscope is the following:

S8007333.jpg

 

This oscillator is around 120 Hz, which should be  the harmonics of 60 Hz from light pollution.

I am not sure exactly why it is stable when the control-loop sign is flipped. This could

be similar to the Pauli trap in the iron trap, because the coil not only provides a force

but also provides the rigidity. The sign of such rigidity depends on the sign of the control

current. If such oscillating rigidity changes at a frequency much higher than the response

frequency of the magnet, it will stablize  the system simply by significantly increasing

the inertial of the magnet.More investigations are essential to completely understand it.

 

For information about Pauli trap, one can look at the wikipedia:

http://en.wikipedia.org/wiki/Quadrupole_ion_trap

 

 

 

  2306   Fri Nov 20 11:14:22 2009 josephb, alexConfigurationComputerstest points working on megatron and we may have filters with switch outputs built in

Alex tooked at the channel definitions (can be seen in tpchn_C1.par), and noticed the rmid was 0. 

However, we had set in testpoint.par the tst system to C-node1 instead of C-node0.  The final number inf that and the rmid need to be equal.   We have changed this, and the test points appear to be working now.

However, the confusing part is in the tst model, the gds_node_id is set to 1.  Apparently, the model starts counting at 1, while the code starts counting at 0, so when you edit the testpoint.par file by hand, you have to subtract one from whatever you set in the model.

In other news, Alex pointed me at a CDS_PARTS.mdl, filters, "IIR FM with controls".  Its a light green module with 2 inputs and 2 outputs.  While the 2nd set of input and outputs look like they connect to ground, they should be iterpreted by the RCG to do the right thing (although Alex wasn't positive it works, it worth trying it and seeing if the 2nd output corresponds to a usable filter on/off switch to connect to the binary I/O to control analog DW.  However, I'm not sure it has the sophistication to wait for a zero crossing or anything like that - at the moment, it just looks like a simple on/off switch based on what filters are on/off.

  2305   Fri Nov 20 11:01:58 2009 josephb, alexConfigurationComputersWhere to find RFM offsets

Alex checked out the old rts (which he is no longer sure how to compile) from CVS to megatron, to the directory:

/home/controls/cds/rts/

In /home/controls/cds/rts/src/include you can find the various h files used.  Similarly, /fe has the c files.

In the h files, you can work out the memory offset by noting the primary offset in iscNetDsc40m.h

A line like suscomms.pCoilDriver.extData[0] determines an offset to look for.

0x108000 (from suscomms )

Then pCoilDriver.extData[#] determines a further offset.

sizeof(extData[0]) = 8240  (for the 40m - you need to watch the ifdefs, we were looking at the wrong structure for awhile, which was much smaller).

DSC_CD_PPY is the structure you need to look in to find the final offset to add to get any particular channel you want to look at.

The number for ETMX is 8, ETMY 9 (this is in extData), so the extData offset from 0x108000 for ETMY should be 9 * 82400.  These numbers (i.e. 8 =ETMX, 9=ETMY) can be found in losLinux.c in /home/controls/cds/rts/src/fe/40m/.  There's a bunch of #ifdef and #endif which define ETMX, ETMY, RMBS, ITM, etc.  You're looking for the offset in those.

So for ETMY LSC channel (which is a double) you add 0x108000 (a hex number) + (9 * 82400 + 24) (not in hex, need to convert) to get the final value of 0x11a160 (in hex).

-----------

A useful program to interact with the RFM network can be found on fb40m.  If you log in and go to:

/usr/install/rfm2g_solaris/vmipci/sw-rfm2g-abc-005/util/diag

you can then run rfm2g_util, give it a 3, then type help.

You can use this to read data.  Just type help read.  We had played around with some offsets and various channels until we were sure we had the offsets right.  For example, we fixed an offset into the ETMY LSC input, and saw the corresponding memory location change to that value.  This utility may also be useful for when we do the RFM test to check the integrity of the ring, as there are some diagnostic options available inside it.

  2304   Fri Nov 20 00:18:45 2009 ranaSummaryCamerasVideo MUX Selection Wiki page

Steve is summarizing the Video Matrix choices into this Wiki page:

http://lhocds.ligo-wa.caltech.edu:8000/40m/Electronics/VideoMUX

Requirements:

Price: < 5k$

Control: RS-232 and Ethernet

Interface: BNC (Composite Video)

Please check into the page on Monday for a final list of choices and add comments to the wiki page.

ELOG V3.1.3-